OpenStudy (gibbs):

MEDAL! 7. at Sydnys birthday party there were 8 pizzas and 15 girls. how much pizza did each girl get if they all got the same amount? 8. mrs. jones has 14 1/4yds. of dress material. if she cuts it into equil pices for each of her 3 daughters, what will be the length of each pice?

OpenStudy (texaschic101):

well...you got 15 girls and you got 8 pizzas....so 15/8 = 1.875 pieces or 1 7/8 pieces per person. next one... (14 1/4) / 3 = 57/4 * 1/3 = 57/12 = 4 3/4 yds

OpenStudy (gibbs):

ty

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

hey @texaschic101 can you help me with a test?

OpenStudy (gibbs):

i can

OpenStudy (anonymous):

ok ill tag you

OpenStudy (gibbs):

ok

OpenStudy (matlee):

Gibbs thats what i wrote on my last answer

OpenStudy (gibbs):

ok sowwy